OpenStudy (anonymous):

Solve the system: 9x^2 + y^2=9 3x-y=3

OpenStudy (anonymous):

first isolate one variable in one of the equations( the easiest) 3x-y = 3 -y=3-3x y = -3 + 3x then fill in y in the other equation: 9x^2 + (-3+3x)^2 = 9 9x^2 + 9 + 9x = 9 9x^2+9x = 0 Here is where I get stuck, but does what I have make any sense to you?

OpenStudy (anonymous):

Ok I figured it out from here thanks!